QUESTION 8 Let x be a binomial random variable with n=5 and p=0.7. Find P(X <= 4). O 0.1681 0.5282 0.4718 0.8319 0.3601
Answer #1

Cation: DE 5, Pa 0.7, Xo Binomial (n=5, P=0.7) P(x4) = P(x=0) + Plx=1) + P(x = 2) +P(X=4) P(x=x) = (2) (3) p (i-emot 3(0-2)*